A researcher wishes to estimate, with 99% confidence, the population proportion of a) No preliminary estimate is available Find the minimum b i F ind the minimum sapi·sue needed using a prior study that found that 32% of the respondents said they think their president c) Compare the results from parts (a) and (b) who think the president of their country can control the price of gasoline Her estimate must be accurate within 4% ofthe tw proportio control the price of gasoline (a) What is the minimum sample size needed assuming that no prior information is available? (Round up to the nearest whole number as needed) (b What is the minin um sample size needed using a prior study mat found that 32% or therespondents aid they tma mer president an a mile po aple Round up to the nearest whole number as needed) (c) How do the results rom (a) and (b) compare A. Having an estimate of the population proportion has no effect on the minimum sample size needed B. Having an estimate o, the population proportion raises the mnmm sample see needed OC. Having an estimate of the population proportion reduces the minimum sample size needed
Answer #1

SOLUTION

\mathbf{given:}\;Margin\;of\;error\;(E)=4\%=0.04

\mathbf{formula:}\;Margin\;of\;error\;(E)=z_{\frac{\alpha }{2}}*\left \{ \sqrt{\frac{p*(1-p)}{n}} \right \}

\mathbf{for \;99\%\; confidence \;level,\; z_{\frac{\alpha }{2}}=2.576}

{\color{Blue} \boldsymbol{(a)}}\;\;p=\mathbf{0.5}\;\mathbf{(no\;prior\;estimate)}

0.04=2.576*\left \{ \sqrt{\frac{0.5*(1-0.5)}{n}} \right \}

\frac{0.04}{2.576}=\sqrt{\frac{0.5*(1-0.5)}{n}}

\frac{0.04}{2.576}=\sqrt{\frac{0.25}{n}}

Sqaure\;on\;both\;sides

\left (\frac{0.04}{2.576} \right )^{2}=\frac{0.25}{n}

n=\frac{0.25}{\left (\frac{0.04}{2.576} \right )^{2}}

n=1036.8

n\approx \mathbf{{\color{Red} 1037}}

--------------------------------------------------------------------------------------------------------------------------------------

{\color{Blue} \boldsymbol{(b)}}\;\;p=\mathbf{0.32}

0.04=2.576*\left \{ \sqrt{\frac{0.32*(1-0.32)}{n}} \right \}

\frac{0.04}{2.576}=\sqrt{\frac{0.32*(1-0.32)}{n}}

\frac{0.04}{2.576}=\sqrt{\frac{0.2176}{n}}

Sqaure\;on\;both\;sides

\left (\frac{0.04}{2.576} \right )^{2}=\frac{0.2176}{n}

n=\frac{0.2176}{\left (\frac{0.04}{2.576} \right )^{2}}

n=902.5

n\approx \mathbf{{\color{Red} 903}}

--------------------------------------------------------------------------------------------------------------------------------------

{\color{Blue} \boldsymbol{(c)}}\;\;{\color{Red} \mathbf{C.}}Having an estimate of the population proportion reduces the minimum sample size needed.